OpenStudy (anonymous):

At a town fair, ten tickets cost eight dollars. How many tickets could you expect to receive for twenty dollars, assuming there is no discount for buying more tickets?

OpenStudy (mathmate):

Use proportions \( \huge \frac{10\ tickets}{ $8 } = \frac{x\ tickets}{$20} \\ then\\ \huge x\ tickets =10\ tickets *\frac{$20}{$8} \) Can you solve for x?